# TraveLER: A Multi-LMM Agent Framework for Video Question-Answering

### v1

@@ Abstract
Recently, Large Multimodal Models (LMMs) have made significant progress in video question-answering using a frame-wise approach by leveraging large-scale, image-based pretraining in a zero-shot manner. While image-based methods for videos have shown impressive performance, a current limitation is that they often overlook how key timestamps are selected and cannot adjust when incorrect timestamps are identified. Moreover, they are unable to extract details relevant to the question, instead providing general descriptions of the frame. To overcome this, we design a multi-LMM agent framework that travels along the video, iteratively collecting relevant information from keyframes through interactive question-asking until there is sufficient information to answer the question. Specifically, we propose TraveLER, a model that can create a plan to “Traverse” through the video, ask questions about individual frames to “Locate” and store key information, and then “Evaluate” if there is enough information to answer the question. Finally, if there is not enough information, our method is able to “Replan” based on its collected knowledge. Through extensive experiments, we find that the proposed TraveLER approach improves performance on several video question-answering benchmarks, such as NExT-QA, STAR, and Perception Test, without the need to fine-tune on specific datasets.
@@ 1 Introduction
Over the last few years, Large Multimodal Models (LMMs) have demonstrated tremendous progress in the area of video understanding, particularly for the video question-answering (VideoQA) domain (Fu et al., 2022; Wang et al., 2022a). More recently, LMMs have been able to achieve impressive results through video-based models (Lin et al., 2023a; Sun et al., 2023; Ye et al., 2022; Li et al., 2022). However, video models require a high level of computational complexity to fine-tune, and annotations are difficult and expensive to collect. As a result, many recent approaches (Xue et al., 2023; Yu et al., 2023; Zhang et al., 2023a) operate on a frame level, leveraging large-scale image-based pretraining in a zero-shot setting.
Despite the effectiveness of image-based LMMs for image tasks, applying them to VideoQA is challenging since using all frames results in high computational demands and redundancy. Thus, many works try to select subsets of frames, either through uniform sampling (Zhang et al., 2023a) or keyframe selection (Yu et al., 2023). However, uniform sampling may skip important information, while keyframe selection methods might select the wrong frames and mislead the model. To address this, we introduce a novel video traversal approach using a “Planner,” an LLM agent that creates a plan to find and extract key information.
Next, we wish to ensure we can capture correct and detailed information from frames when executing the plan. Yet, the common captioning approach provides general descriptions for the frame, whereas answering questions often requires more specific details. Moreover, not all elements of the frame are relevant to the question and may even be misleading. As a result, we propose an interactive question-answering process using a “Locator” to locate and extract the most relevant and fine-grained details from each frame. Specifically, we use two LMM agents, one who asks questions about the frame and another who answers them.
Nevertheless, it can be difficult for the model to collect all necessary details in a single pass, and extracting incorrect information may be misleading. Hence, we introduce an iterative approach using an ”Evaluator” that reviews collected information after each iteration and evaluates if it has enough information to answer the question. If there is, the answer is selected; otherwise, the new information is used to ”Replan” and begin a new iteration.
@@
Consider the example in Figure 1. Suppose we are asked ”why the boy turned over in the middle of the video”. In the first iteration, our method uses temporal cues from the question to skip to the middle of the video and asks questions to find the relevant frames. In the next iteration, we gather more information. Asking about what the boy is doing, we learn that he is ”standing up at the bottom of the slide” and is not looking at anything specific, which informs us that the boy is no longer ”sitting down” (choice B) or ”resting on the yellow object” (choice D). To eliminate these choices, we must confirm that the boy does not sit back down again by traveling to a timestamp near the end of the video. Finally, since we have collected enough information and followed the plans, we can select the right choice that the baby turns over to be on his stomach ”to get down on slide” (Choice E).
@@
Our proposed approach – Traverse, Locate, Evaluate, and Replan (TraveLER), is a modular, multi-LMM agent framework for video question answering. Our framework is composed of four main stages, each with LLM or LMM “agents” that interact with each other through the different stages. First, in the Traversal stage (“traverse”), an agent creates a plan to answer the question. In the Locator stage (“locate”), an agent uses the plan to decide which timestamp of the video to select. The corresponding frames are then sent to another agent, which asks questions and stores the answers in a memory bank for future iterations. Finally, in the Evaluator stage (“evaluate”), an agent reviews all collected information and decides whether to answer or create a modified plan (“replan”) to start the next iteration if necessary.
@@
To summarize, our main contributions are as follows: (i) We introduce TraveLER, a modular multi-LMM agent framework for video question-answering. (ii) Our proposed TraveLER method does not require task-specific fine-tuning or video annotations, as well as being easy to employ with several different LLM or LMMs. (iii) Our method shows improved performance on multiple difficult video question-answering benchmarks such as NExT-QA, Perception Test, and STAR, highlighting the effectiveness of our approach.
@@ 2 Related Work
Video Question-Answering. Video question-answering (VideoQA) involves answering free-form or multiple choice questions given an input video. In comparison to image question answering, VideoQA poses unique challenges because it often requires strong temporal understanding and the ability to deal with long input sequences. Many recent works have focused on training end-to-end video-language models (Fu et al., 2022; Sun et al., 2023; Wang et al., 2022a; Ye et al., 2022; Yu et al., 2022; Li et al., 2022), but doing so remains challenging due to computational constraints and difficulties in architecture scaling. As a result, many approaches adopt pretrained image models to the video domain by extracting information independently from each frame (Xue et al., 2023; Yu et al., 2023; Zhang et al., 2023a). In this work, we design a framework that builds an adaptive plan to traverse through the video to identify keyframes and extract relevant information using a question-answering approach.
@@
Large Multimodal Models for Video Understanding. LMMs have been shown to be extremely useful for VideoQA. Some methods use supervised or contrastive training to perform video-LMM pretraining (Zhao et al., 2022; Yang et al., 2023; Chen et al., 2023b), while others adapt existing LMMs and use instruction tuning to adapt them to the video domain (Zhang et al., 2023c; Maaz et al., 2023; Lin et al., 2023a). However, recent improvements in LMM capabilities have allowed for many strong approaches for few-shot (Alayrac et al., 2022; Wang et al., 2022b) and zero-shot VideoQA (Yang et al., 2022; Wang et al., 2022b). In particular, zero-shot methods such as LLoVi (Zhang et al., 2023a), use pre-trained LMMs to generate captions for each frame in the video. Nevertheless, uniformly sampling frames at random may result in the model missing important visual information and focusing on unimportant frames  (Wu et al., 2019; Lei et al., 2021b). Recent works such as SeViLA (Yu et al., 2023) addressed this problem by performing parameter-efficient finetuning using captions to identify keyframes (Lu et al., 2022; Buch et al., 2022; Qian et al., 2023), but this requires fine-tuning on specific datasets. In contrast to these works, which select all keyframes in a single pass, we introduce a zero-shot, iterative method that repeatedly gathers data from various timestamps until enough information is collected to correctly answer the question.
@@
LMM-based Agents for Videos. The strong reasoning abilities of LLMs  (Brown et al., 2020; Chung et al., 2022) have made them effective in LLM-based agent approaches for videos, where an LLM performs much of the reasoning after collecting information from different modules (Chen et al., 2023a; Lin et al., 2023b; Zhang et al., 2023b; Zeng et al., 2022). For example, Socratic Models (Zeng et al., 2022) proposes a method to reason about videos based on generated audio transcriptions and CLIP frame similarity scores, while other works like VideoChatCaptioner (Chen et al., 2023a) proposes a way to caption videos through chat dialogues between an LLM and a LMM. Unlike these works, our method utilizes a novel video traversal approach and an iterative, planning-based information gathering process.
@@ 3 TraveLER Framework
To design a robust approach that can find the correct keyframes and extract the most relevant information for VideoQA, we propose a modular LMM agent framework that Traverses, Locates, Evaluates, and Replans iteratively (TraveLER). We begin by describing the general LLM and LMM architectures (Section 3.1), then introduce each component of our pipeline (Section 3.2), and implementation details (Section 3.3). Our method is illustrated in Figure 2.
@@ 3.1 Preliminaries
Large language and multimodal model architectures. LLMs are text-conditioned generative models. Given a prompt $P_{\text{in}}$, they encode it into a fixed language embedding $l$ in an embedding space $f(\cdot)$ and then use this to produce a text response $R$: $R=f(l(P_{\text{in}}))$.
Similarly, Large Multimodal Models (LMMs) are adapted to jointly reason over vision and language modalities. To map the different modalities into the shared embedding space $f(\cdot)$, the image $I$ is encoded using a trainable encoder $v$ and the prompt $P_{\text{in}}$ is encoded using a fixed language embedding $l$. The LMM outputs a text response $R$: $R=f(v(I),l(P_{\text{in}}))$.
Video question-answering. VideoQA involves viewing a video and answering questions. The model is usually evaluated through top-1 accuracy, in which it chooses the best answer out of a set of possible choices. Specifically, given a question $Q$, video input $V$ consisting of a set of frames $\{f_{1},\cdots,f_{n}\}$, and set of choices $C=\{c_{1},\cdots,c_{n}\}$, the model is asked to choose the best $c_{i}$ to answer $Q$. Next, we introduce each component of our method.
@@ 3.2 TraveLER Components
Our method uses a memory bank $M$ to store collected information, which allows information to persist and to be updated as we proceed through different iterations. We initialize $M$ with captions of 5 evenly sampled frames throughout the video. We find that this memory initialization gives the model good context for the general idea of the video, and performs better than starting with an empty memory $M$. After the first iteration, we add information iteratively using the Extractor module, which will be discussed later.
Next, we input the prompt $P_{\mathrm{T}}^{(1)}$ into $\text{LLM}_{\text{planner}}$, which returns response $R_{T}$, a step-by-step plan on how to traverse through the video and what information the model needs to collect.
@@
Our next step is to use the plan $R_{T}$ in the Locator stage to locate keyframes and extract the information that we will use to answer the question.
Locator. The Locator is a component that consists of two submodules, the Retriever and the Extractor. The Retriever selects the timestamps of the next frames to view, while the Extractor extracts relevant information from these frames, using a question-answering process. Next, we discuss each component in more detail.
(i) Retriever: The Retriever carries out the given plan $R_{T}$ by selecting which frames to view next. The Retriever is an LLM-based submodule whose goal is to use the collected information $M$ to find the next timestamp $t$ to select in order to fulfill the plan $R_{T}$. The task prompt $P_{R}$ is an instruction that contains information about the video length and asks which timestamp to view next. Thus, we insert the question $Q$, plan $R_{T}$, and collected information $M$ into the task prompt $P_{R}$ to create the new prompt $P_{\mathrm{R}}^{(1)}$: $P_{\mathrm{R}}^{(1)}=``[P_{R}][Q][R_{T}][M]"$.
Given prompt $P_{\mathrm{R}}^{(1)}$, the LLM in the Retriever, $\text{LLM}_{\text{retriever}}$, returns $t$, or the next set of timestamps. The module then retrieves frames $I_{t}$ at timestamp $t$.
(ii) Extractor: The Extractor is a significant part of our method because it allows us to capture more relevant and question specific details from the visual input, unlike using only captions. We pass the frames selected by the retriever $I_{t}$ into the Extractor submodule, which consists of two large models: an LLM, $\text{LLM}_{\text{extractor}}$, to generate context-dependent questions about the frames $I_{t}$, and a different vision-language LMM, $\text{LMM}_{\text{extractor}}$, whose job is to extract the desired information from the same frames.
In this module, we first generate a general caption $c_{t}$ for frame $I_{t}$ using the $\text{LMM}_{\text{extractor}}$. Then, we concatenate the caption $c_{t}$, plan $R_{T}$, and memory $M$, and the Extractor task prompt $P_{E}$, which is an instruction that asks to use available information to create 3 questions to ask about the current frame. This results in the new prompt $P_{\mathrm{E}}^{(1)}$: $P_{\mathrm{E}}^{(1)}=``[P_{E}][c_{t}][R_{T}][M]"$.
@@
$\{a_{1},a_{2},\cdots,a_{n}\}=\text{LMM}_{\text{extractor}}(v(I_{t}),l(\{q_{1},q_{2},\cdots,q_{n}\}))$
@@
Finally, to use this collected information in future iterations, we update our memory bank $M$. To do this, we use the timestamp $t$ of $I_{t}$ as our key and the question-answer pair list as the value, and append this to our memory $M$.
If the memory bank dictionary $M$ is too long, we summarize it by using the memory bank as an input to another LLM and instruct it to make the memory bank entries more concise, while retaining the same keys and format. This output becomes our new memory bank.
@@
We use this prompt $P_{\mathrm{V}}^{(1)}$ as input into the LLM in the Evaluator, $\text{LLM}_{\text{evaluator}}$, which evaluates if there is enough information to answer the question and if the plan has been completely followed. If both are true, $\text{LLM}_{\text{evaluator}}$ outputs the best choice $c*$ to answer the question $Q$. Otherwise, it provides an explanation $E$ on why there is not enough information and gives this explanation to the Planner to start a new iteration of the process.
Re-planning. After each iteration, if the evaluator decides that there is not enough information to answer the question $Q$ or if the plan $P$ has not been completed, the existing memory $M$ will be provided to the Planner in the next iteration, in addition to an explanation $E$ for why an answer was not chosen. The Planner then outputs a new plan, restarting the process. We also implement a limit on the number of iterations a question can take, to prevent infinite loops. After this limit is reached, we force the Evaluator to choose the best choice.
3.3 Implementation Details
Here, we discuss how we implement various components of our framework. The code will be released upon acceptance. More implementation details, such as prompts and dataset-specific details are in the Supplementary in Section B.
Agent model selection. Our modular approach has the benefit of allowing us to easily swap in different LLMs and LMMs (see Section 4.4). For our main experiments, we use LLaVA-1.6 for $\text{LMM}_{\text{extractor}}$ and GPT-4 for $\text{LLM}_{\text{planner}}$, $\text{LLM}_{\text{retriever}}$, $\text{LLM}_{\text{extractor}}$, and $\text{LLM}_{\text{evaluator}}$.
Multi-frame selection. We also allow for the Retriever to select multiple frames instead of a single frame. This helps the model to better capture events that happen quickly or require more context to recognize. For example, if we want to find the action of a ”a woman clapping her hands”, single frame selection may cause us to incorrectly assume the woman is not clapping if we view the frame where their hands are apart. We do this by creating an optional parameter called window size. The window size refers to the number of frames the Retriever extracts each time. When the window size is non-zero, the Retriever still specifies a single timestamp to go to, but when retrieving the frame at that timestamp we take the number of frames specified by the window size before and after the selected frame as well.
Table 1: Zero-shot (ZS) results on NExT-QA. For fair comparisons, we gray out methods with fine-tuned components in their model. The best scores are in bold.
Model Cau. (%) Tem. (%) Des. (%) Avg. (%)
MC-ViT-B (Balažević et al., 2024) - - - 60.6
SeViLA (Yu et al., 2023) 61.3 61.5 75.6 63.6
MC-ViT-L (Balažević et al., 2024) - - - 65.0
InternVideo (Wang et al., 2022a) 43.4 48.0 65.1 49.1
VFC (Momeni et al., 2023) 45.4 51.6 64.1 51.5
ViperGPT (Surís et al., 2023) - - - 60.0
BLIP-2concat (Yu et al., 2023) 59.7 60.8 73.8 62.4
BLIP-2voting (Yu et al., 2023) 59.1 61.3 74.9 62.7
ProViQ (Choudhury et al., 2023) - - - 63.8
LLoVi (Zhang et al., 2023a) 69.5 61.0 75.6 67.7
TraveLER (ours) 70.0 (+0.5) 60.5 78.2 (+2.6) 68.2 (+0.5)
Table 2: ZS results on STAR.
Model Int. Seq. Pre. Fea. Avg.
SeViLA 48.3 45.0 44.4 40.8 44.6
Flamingo-9B - - - - 41.8
InternVideo 43.8 43.2 42.3 37.4 41.6
BLIP-2voting 41.8 39.7 40.2 39.5 40.3
BLIP-2concat 45.5 41.8 41.8 40.0 42.2
TraveLER (ours) 43.3 45.9 45.8 42.8 44.9  (+2.7)
Table 3: ZS results on Perception Test.
Model Test Accuracy
@@ 3.3 Implementation Details
TraveLER (ours) 50.2  (+4.5)
4 Experiments and Results

### v2

# TraveLER: A Modular Multi-LMM Agent Framework for Video Question-Answering

@@ Abstract
Recently, image-based Large Multimodal Models (LMMs) have made significant progress in video question-answering (VideoQA) using a frame-wise approach by leveraging large-scale pretraining in a zero-shot manner. Nevertheless, these models need to be capable of finding relevant information, extracting it, and answering the question simultaneously. Currently, existing methods perform all of these steps in a single pass without being able to adapt if insufficient or incorrect information is collected. To overcome this, we introduce a modular multi-LMM agent framework based on several agents with different roles, instructed by a Planner agent that updates its instructions using shared feedback from the other agents. Specifically, we propose TraveLER, a method that can create a plan to “Traverse” through the video, ask questions about individual frames to “Locate” and store key information, and then “Evaluate” if there is enough information to answer the question. Finally, if there is not enough information, our method is able to “Replan” based on its collected knowledge. Through extensive experiments, we find that the proposed TraveLER approach improves performance on several VideoQA benchmarks without the need to fine-tune on specific datasets. Our code is available at https://github.com/traveler-framework/TraveLER.
TraveLER: A Modular Multi-LMM Agent Framework
for Video Question-Answering
[…]
@@ 1 Introduction
Over the last few years, Large Multimodal Models (LMMs) have demonstrated tremendous progress in the area of video understanding, particularly for the video question-answering (VideoQA) domain (Fu et al., 2023; Wang et al., 2024). More recently, LMMs have been able to achieve impressive results through video-based models (Lin et al., 2023a; Sun et al., 2022; Ye et al., 2023; Li et al., 2023). However, video models require a high level of computational complexity to fine-tune, and annotations are difficult and expensive to collect. As a result, many recent approaches (Xue et al., 2023; Yu et al., 2023; Zhang et al., 2023a) operate on the frame level, leveraging large-scale image-based pretrained models in a zero-shot manner.
Moreover, these models may need to do several tasks simultaneously in a single step. In particular, they should identify the correct events in videos by understanding what information is relevant and ignoring irrelevant information. Next, they would need to extract specific and question-relevant visual details and use them to answer the question. As such, one iteration might not be enough to collect all the necessary information. For example, many current approaches use simple captioning, which is often too general to extract specific details, or they might miss important events. In these cases, current approaches cannot revisit the video to find additional information. Decomposing this process into different components allows each component to adapt to newly collected information.
To address this, we introduce a modular multi-LMM agent framework for VideoQA. A Planner agent instructs agents in charge of different tasks, such as navigating through the video, extracting visual information through a question-answering process, and reviewing this information to select an answer. Feedback from the agents is then passed on to each other and back to the Planner, who uses the feedback to update its instructions. In this way, we can revisit the video to narrow our focus or expand our search to more relevant information and extract specific details to answer the question.
@@
Consider the example in Figure 1. Suppose we are asked why the boy turned over in the middle of the video. In the first iteration, our method uses temporal cues from the question to skip to the middle of the video and asks questions to find the relevant frames. In the next iteration, we gather more information. Asking about what the boy is doing, we learn that he is “standing up at the bottom of the slide" and is not looking at anything specific, which informs us that the boy is no longer “sitting down" (choice B) or “resting on the yellow object" (choice D). To eliminate these choices, we must confirm that the boy does not sit back down again by traveling to a timestamp near the end of the video. Finally, since we have collected enough information and followed the plans, we can select the right choice that the boy turns over to be on his stomach “to get down from slide” (Choice E).
@@
Our proposed approach – Traverse, Locate, Evaluate, and Replan (TraveLER), has four main stages. First, in the Traversal stage, an agent creates a plan to answer the question. In the Location stage, an agent uses the plan to decide which timestamp of the video to select. The corresponding frames are then sent to another agent, which asks questions and stores the answers in a memory bank for future iterations. Finally, in the Evaluation stage, an agent reviews all collected information and decides whether to answer or create a modified plan (Replan) to start the next iteration if necessary.
@@
To summarize, our main contributions are: (i) We introduce TraveLER, a modular multi-LMM agent framework for video question-answering. (ii) Our method shows improved performance on multiple difficult video question-answering benchmarks, such as NExT-QA, EgoSchema, Perception Test, and STAR. (iii) Our method is easy to employ with different LLMs and LMMs, highlighting the effectiveness of our modular approach.
@@ 2 Related Work
Video Question-Answering. VideoQA involves answering free-form or multiple-choice questions given an input video. Compared to image question answering, VideoQA poses unique challenges because it often requires strong temporal understanding and the ability to deal with long input sequences. Many recent works have focused on training end-to-end video-language models (Fu et al., 2023; Sun et al., 2022; Wang et al., 2024; Ye et al., 2023; Yu et al., 2022; Li et al., 2023), but doing so remains challenging due to computational constraints and difficulties in architecture scaling. As a result, many approaches adopt pretrained image models to the video domain by extracting information independently from each frame (Xue et al., 2023; Yu et al., 2023; Zhang et al., 2023a). Here, we design a framework that builds an adaptive plan to find and extract relevant information using a question-answering approach.
@@
LMMs for Video Understanding. LMMs have been shown to be extremely useful for VideoQA. Some methods use supervised or contrastive training to perform video-LMM pretraining (Zhao et al., 2023; Yang et al., 2023; Chen et al., 2024), while others adapt existing LMMs and use instruction tuning to adapt them to the video domain (Zhang et al., 2023b; Maaz et al., 2024; Lin et al., 2023a). However, recent improvements in LMM capabilities have allowed for many strong approaches for few-shot (Alayrac et al., 2022; Wang et al., 2022) and zero-shot VideoQA (Yang et al., 2022). In particular, zero-shot methods, such as LLoVi (Zhang et al., 2023a), use pre-trained LMMs to generate captions for each frame in the video. Nevertheless, uniformly sampling frames at random may miss important visual information and focus on unimportant frames  (Wu et al., 2019; Lei et al., 2021b). Recent works like SeViLA (Yu et al., 2023) addressed this problem by performing parameter-efficient finetuning using captions to identify keyframes (Lu et al., 2022; Buch et al., 2022; Qian et al., 2023), but this requires fine-tuning on specific datasets. Unlike these works, which select all keyframes in a single pass, we introduce a novel iterative and modular approach instructed by a planner.
@@
LMM-based Agents for Videos. The strong reasoning abilities of LLMs  (Brown et al., 2020; Chung et al., 2022) have made them effective in LLM-based agent approaches for videos, where an LLM performs much of the reasoning after collecting information from different modules (Chen et al., 2023; Lin et al., 2023b; Zhang et al., 2024; Zeng et al., 2023). For example, Socratic Models (Zeng et al., 2023) proposes a method to reason about videos based on generated audio transcriptions and CLIP frame similarity scores, while other works like VideoChatCaptioner (Chen et al., 2023) proposes a way to caption videos through chat dialogues between an LLM and a LMM. Recently, there have also been works that use program generation using an LLM to answer questions  (Choudhury et al., 2023; Min et al., 2024). However, these works still use a single-pass approach and provide very general captions. In contrast, our work uses an iterative question-answering process to extract specific, relevant information in the frame.
@@ 3 TraveLER Framework
We begin by describing some background on the LLM and LMM architectures (Section 3.1), then introduce each component of our framework (Section 3.2) and implementation details (Section 3.3). Our method is shown in Figure 2.
@@ 3.1 Preliminaries
Large language and multimodal models. LLMs are text-conditioned generative models. Given a prompt $P$, they encode it into a fixed language embedding $l$ in an embedding space $f(\cdot)$ and use this to produce text response $R$: $R=f(l(P))$. Similarly, Large Multimodal Models (LMMs) are adapted to jointly reason over vision and language. To map different modalities into the shared embedding space $f(\cdot)$, an image $I$ is encoded using an encoder $v$, and the prompt $P$ is encoded using a fixed language embedding $l$. The LMM outputs a textual response $R$: $R=f(v(I),l(P))$.
Video question-answering. VideoQA involves viewing a video and answering questions. The model is usually evaluated through top-1 accuracy, which chooses the best answer out of a set of possible choices. Specifically, given a question $Q$, video input $V$ consisting of a set of frames $\{I_{1},\cdots,I_{n}\}$, and a set of choices $C=\{c_{1},\cdots,c_{n}\}$, the model is asked to choose the best $c_{i}$ to answer $Q$. Next, we introduce each component of our method.
@@ 3.2 TraveLER Components
Our method uses a memory bank $M$ to store collected information, which allows information to persist and to be updated as we proceed through different iterations. We initialize $M$ with captions of 5 evenly sampled frames throughout the video. We find that this memory initialization gives the model good context for the general idea of the video, and performs better than starting with an empty memory $M$. After the first iteration, we add information iteratively using the Extractor module, which we discuss later in this section.
Next, we input the prompt $P_{\mathrm{T}}^{(1)}$ into $\text{LLM}_{\text{planner}}$, which returns response $R_{T}$, a step-by-step plan on how to traverse through the video and what information should be collected. This plan is revised in future iterations using collected information.
@@
Our next step is to use the plan $R_{T}$ in the Locator stage to locate relevant events and extract the information that we will use to answer the question.
Locator. The Locator is a component that consists of two submodules, the Retriever and the Extractor. The Retriever selects the timestamps of the next frames to view, while the Extractor extracts relevant information from these frames using a question-answering process. Next, we discuss each component in more detail.
(i) Retriever: The Retriever carries out the given plan $R_{T}$ by selecting which frames to view next. The Retriever is an LLM-based submodule that finds the next timestamp $t$ to view, given the plan $R_{T}$, collected information $M$, and video metadata (frame rate, length). The task prompt $P_{R}$ is an instruction that contains information about the video length and asks which timestamp to view next. Thus, we insert the question $Q$, the plan $R_{T}$, and the collected information $M$ into the task prompt $P_{R}$ to create the new prompt $P_{\mathrm{R}}^{(1)}$: $P_{\mathrm{R}}^{(1)}=``[P_{R}][Q][R_{T}][M]"$.
Given prompt $P_{\mathrm{R}}^{(1)}$, the LLM in the Retriever, $\text{LLM}_{\text{retriever}}$, returns $t$, the next set of timestamps. Then, it retrieves frames $I_{t}$ at timestamp $t$.
(ii) Extractor: The Extractor is important because it allows us to capture more relevant and question-specific details from the visual input, unlike using only captions. We pass the frames selected by the retriever $I_{t}$ into the Extractor submodule, which consists of two large models: $\text{LLM}_{\text{extractor}}$, to generate context-dependent questions about the frames $I_{t}$, and a different vision-language model $\text{LMM}_{\text{extractor}}$, which extracts the desired information from the same frames. We note that we use both an LLM and LMM since the LLM is better at reasoning about the plan and the collected information, while the LMM is able to collect visual information requested by the LLM.
We first generate a general caption $c_{t}$ for frame $I_{t}$ using the $\text{LMM}_{\text{extractor}}$. Then, we concatenate the caption $c_{t}$, plan $R_{T}$, and memory $M$, and the Extractor task prompt $P_{E}$, which is an instruction that asks to use available information to create 3 questions to ask about the current frame. This results in the new prompt $P_{\mathrm{E}}^{(1)}$: $P_{\mathrm{E}}^{(1)}=``[P_{E}][c_{t}][R_{T}][M]"$.
@@
$\{a_{1},\cdots,a_{n}\}=\text{LMM}_{\text{extractor}}(v(I_{t}),l(\{q_{1},\cdots,q_{n}\}))$
@@
Finally, to use this collected information in future iterations, we update our memory bank $M$. To do this, we use the timestamp $t$ of $I_{t}$ as our key and the question-answer pair list as the value, and append this to our memory $M$. If the memory bank dictionary $M$ is too long, we summarize it by using the memory bank as an input to another LLM and instruct it to make the memory bank entries more concise, while retaining the same keys and format. This output becomes our new memory bank.
@@
We use this prompt $P_{\mathrm{V}}^{(1)}$ as input into the LLM in the Evaluator, $\text{LLM}_{\text{evaluator}}$, which evaluates if there is enough information to answer the question and if the plan has been completely followed. If both are true, $\text{LLM}_{\text{evaluator}}$ outputs the best choice $c^{*}$ to answer the question $Q$. Otherwise, it provides an explanation $E$ on why there is insufficient information and gives this explanation to the Planner to start a new iteration of the process.
Re-planning. After each iteration, if the evaluator decides that there is not enough information to answer the question $Q$ or if the plan $P$ has not been completed, the existing memory $M$ will be provided to the Planner in the next iteration, in addition to an explanation $E$ for why an answer was not chosen. The Planner then outputs a new plan, restarting the process. We also set a limit on the number of iterations a question can take to prevent infinite loops. After reaching this limit, we force the Evaluator to choose the best choice.
Summarizer. The Summarizer is an optional module used for some datasets to summarize the information. It is given the question $Q$, choices $C$, memory bank $M$, and task prompt $P_{S}$, which is an instruction to summarize the information in each timestamp of the memory bank $M$. The outputted summary is then used to replace the original memory bank $M$. This process helps reduce the amount of information passed to the LLM in future steps, which may struggle with very long inputs.
Table 1: Results on Datasets. We show zero-shot results on different datasets. For fair comparisons, we gray out methods with fine-tuned components in their model.
(a) NExT-QA
Model Cau. Tem. Des. Avg.
SeViLA 61.3 61.5 75.6 63.6
MC-ViT-L - - - 65.0
InternVideo 43.4 48.0 65.1 49.1
ViperGPT - - - 60.0
ProViQ - - - 63.8
LLoVi 69.5 61.0 75.6 67.7
TraveLER 70.0 60.5 78.2 68.2 (+0.5)
(b) EgoSchema (Full)
Model Acc.
mPLUG-Owl 31.1
InternVideo 32.1
LongViViT 33.3
Vamos 48.3
LLoVi 50.3
TraveLER 53.3 (+3.0)
(c) STAR
Model Avg.
SeViLA 44.6
Flamingo-9B 41.8
InternVideo 41.6
BLIP-2voting 40.3
BLIP-2concat 42.2
TraveLER 44.9  (+2.7)
(d) Perception Test
Model Acc.
@@
TraveLER 50.2  (+4.5)
3.3 Implementation Details
Here, we discuss how we implement various components of our framework. More implementation details, such as prompts and dataset-specific details are in the Supplementary in Section B.
Agent model selection. Our modular approach has the benefit of allowing us to easily swap in different LLMs and LMMs (see Section 4.4). For our main experiments, we use LLaVA-1.6 (Liu et al., 2023) for $\text{LMM}_{\text{extractor}}$ and GPT-3.5/GPT-4 (OpenAI et al., 2023) for $\text{LLM}_{\text{planner}}$, $\text{LLM}_{\text{retriever}}$, $\text{LLM}_{\text{extractor}}$, and $\text{LLM}_{\text{evaluator}}$.
Multi-frame selection. We also allow the Retriever to select multiple frames instead of one. This helps to capture better events that happen quickly or require more context to recognize. For example, if we want to find the action of "a woman clapping her hands", a single frame selection may cause us to incorrectly assume the woman is not clapping if we view the frame where their hands are apart. We do this by creating an optional parameter called window size. The window size refers to the number of frames the Retriever extracts each time. When the window size is non-zero, the Retriever still specifies a single timestamp to go to, but when retrieving the frame at that timestamp, we also take the number of frames specified by the window size before and after the selected frame.
4 Evaluation